John want to build a box that will hold 400ft. cubed of salt. If it has a square base of 5ft x 5ft, how high must the box be?
I tried: 400ft cubed equals 5 squared but I don't know how to complete it!

Answer by Alan3354(69443)

John want to build a box that will hold 400ft. cubed of salt. If it has a square base of 5ft x 5ft, how high must the box be?
------------------------------
Vol = l*w*h
400 = 5*5*h = 25h
h = 400/25
h = 16 feet